Operation Summary for June 27, 2011 - June 29, 2011 |
Monday June 27 The day shift began with Linac down due to an interlock problem and then for safety system tests, with Operations monitoring the Tevatron (TEV) store 8844 (221.65E30), the Recycler stash (127.8E10), and the Antiproton Source (Pbar) stack (24.3E10), with the Main Injector (MI), Booster, and the NuMI, MiniBooNE, and Meson FTBF T-978 experiments without beam due to Linac being down, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Linac technicians began working on two RF stations (LRF3 & 4). Operators prepared MTA for an access at 9:06 AM. Interlock personnel completed their Linac safety system testing at 9:08 AM. Linac RF work completed at 10:39 AM. At 3:40 PM, two Linac Klystron RF stations (KRF3 & 6) tripped off on a “Reflected Power” indication; they reset. During the evening shift, at 4:21 PM, BRF5 tripped off on a “Driver” fault and would not reset; Booster RF personnel brought it back online. Operations disabled beam to T-978 at 6:01 PM. Operations terminated store 8844 at 9:40 PM, put the TEV into HEP shot setup, and established store 8845 at 11:04 PM with an average initial luminosity (AIL) of 389.78E30. Operators reported that there was a MI Power Supply interlock bypass loop trip during the last transfer of antiprotons from the Pbar Accumulator to the TEV; that transfer was lost. Tuesday June 28 The midnight shift began with Operations monitoring store 8845 (282.86E30) and the stash (40.4E10), with Pbar stacking (25.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 12:34 AM, Operations received a high water alarm for the Fixed Target Neutrino beam line NM2 retention pit. Operators notified an ES&H officer and then accessed the area. They reported, at 12:58 AM, that the water was 10” below the top of the pit. Operators reported finding a small leak on NM2V and valved its LCW off; another magnet, NM2Q1, was also on the same LCW line as NM2V. Operators turned off power to NM2V, NM2Q1, and NM2Q2. There was no beam to this Fixed Target area. BRF6 tripped off at 1:08 AM; it reset. At 2:08 AM, BRF19 and 16 tripped off; BRF19 reset. BRF16 was sparking a lot and causing all the other Booster RF stations to spark. Operators contacted a RF person and then made adjustments on BRF16. Operations established beam to T-978 at 4:24 AM. BRF16 tripped off twice more during the shift; it reset. During the day shift, a Switchyard person reported at 8:39 AM that the Switchyard vacuum was degrading. Operators prepared NM2 for access at 8:52 AM. BRF16 tripped off and resumed sparking at 9:11 AM; operators notified a Booster RF person. BRF16 then went on to trip off six more times in 20 minutes. NuMI personnel began a 30-minute target scan at 9:41 AM. Operations terminated store 8845 at 10:17 AM and then prepared CDF and D0 for accesses. Operators reported at 10:25 AM that cryo system personnel were beginning their flywheel change out on the wet engines at TEV sectors A1 and A2. Interlock technicians began TEV safety system checks at 10:45 AM and on TEV beam monitors in sectors A-E. At 11:04 AM, the NuMI horn tripped off on a relay interlock indication; it eventually reset, but then tripped off twice more; operators informed NuMI personnel. During the evening shift, CDF came out of access at 5 PM. D0 came out of access at 6:30 PM. Operations disabled beam to T-978 at 7:45 PM; this completed T-978’s present run. Operations put the TEV into shot setup and established store 8848 at 11:43 PM with an AIL of 354.32E30. Wednesday June 29 The midnight shift began with Operations monitoring store 8848 (234.65E30) and the stash (175.1E10), with Pbar stacking (16.8E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 2:09 AM, operators reported a minor power glitch that tripped off quite a few power supplies and other devices; everything reset. At 5:22 AM, operators reported that the MI South Anode Power Supply tripped off along with the MI beam permit. Operators investigated and then reset the power supply and permit. Operations turned off the Booster at 7:35 AM and prepared Booster, MI-12, and Pbar for accesses. The Plans for Wednesday The plans for today are to complete the Pbar and Linac safety system tests and for Linac technicians to work on LRF3 & 4, all of these occurring, hopefully, during the Booster access to work on the Booster RF system. Once the Booster RF system work is completed, Operations will reestablish stacking, stashing, and storing. Later this afternoon, after store 8848 is terminated, the Run Coordinator will allow D0 personnel to make a two-hour access. |
|
More Information Comments and Suggestions |
| last modified 06/29/2011 email Fermilab |
| Security, Privacy, Legal | |